OpenStudy (anonymous):

Find the derivative of F(y)= (1/y^2 - 3/y^4)(y+5y^3)

OpenStudy (dumbcow):

\[(\frac{1}{y^{2}}-\frac{3}{y^{4}})(y+5y^{3})\] correct?

OpenStudy (anonymous):

yes!

OpenStudy (dumbcow):

ok first combine fractions \[(\frac{y^{2}-3}{y^{4}})(y+5y^{3})\] multiply \[\frac{5y^{5}-14y^{3}-3y}{y^{4}}\] \[=5y-14y^{-1}-3y^{-3}\] now it should seem easier to take derivative :)

OpenStudy (anonymous):

thank you so much! The algebra gets me every time!

OpenStudy (dumbcow):

no problem